# publications.bib

### v3

@@ 3. Mantovan’s Formula and the Hodge-Tate Period Morphism
The perfectoid Igusa variety $\mathfrak{Ig}^{b}_{C}$ comes equipped with an action of $\mathcal{J}_{b}$. Namely, using [CS1, Proposition 4.2.11], we get an action on the trivialization of $\mathbb{X}_{b}$, as in the moduli description in equation (2). This action extends to the formal model, giving rise to the action of $\mathcal{J}_{b}$ the generic fiber. This allows us to form the map of $v$-stacks:
@@
We would like to say $\pi_{\mathfrak{Ig}}^{b}$ pulls back to the map $\pi^{b}_{\mathrm{HT}}$. However, as seen in Corollary 3.6, we need to account for the additional points in the fiber of the Hodge-Tate period morphism that are not seen by the perfectoid Igusa varieties $\mathfrak{Ig}^{b}_{C}$. To capture this, we need to show that $\pi_{\mathfrak{Ig}}^{b}$ also extends to the partial minimal compactification. We have the following.
@@
[Zha, Corollary 9.43] Assuming 1.10, the action of $\mathcal{J}_{b}$ on the perfectoid Igusa variety $\mathfrak{Ig}^{b}_{C}$ extends uniquely to an action on $\mathfrak{Ig}^{b,*}_{C}$. In particular, by functorality of the formation of the canonical compactification ([Ecod, Proposition 18.6]), we have a map
@@
extending $\pi_{\mathfrak{Ig}}^{b}$. This action preserves the boundary $\partial\mathfrak{Ig}^{b,*}$, so, in particular, we also get a map
$\pi_{\mathfrak{Ig}}^{b,\partial}:[(\overline{\mathfrak{Ig}}^{b,*}_{C}\setminus\overline{\partial\mathfrak{Ig}}^{b,*}_{C})/\mathcal{J}_{b}]\rightarrow[\mathop{\rm Spd}(C)/\mathcal{J}_{b}]$
by restriction.
@@
as explained in the proof of Proposition 3.7. Under Assumption 1.10, we can apply the algebraic form of Hartogs’ principle (See for example [Schtor, Proposition III.2.9]) to the open inclusion $g_{b,m}$ to conclude an isomorphism of global sections via restriction. This gives the corresponding claim for the map $g_{b}$ of perfect schemes. In particular, we have an isomorphism
@@
on global sections. Taking generic fibers of the corresponding integral models, we claim that we obtain an isomorphism
(6) ${}\mathcal{O}(\mathfrak{Ig}^{b}_{C})\simeq\mathcal{O}(\mathfrak{Ig}^{b,*}_{C}),$
of global sections. We need to be a bit careful here with analytic sheafification. In particular, for an index set $I$, we let $\{U_{i}\}_{i\in I}$ be an affine covering of $\mathrm{Ig}^{b}$, and compute global sections via the Čech complex
(7) ${}0\rightarrow\mathcal{O}(\mathrm{Ig}^{b})\rightarrow\prod_{i\in I}\mathcal{O}(U_{i})\rightarrow\prod_{i,j\in I}\mathcal{O}(U_{i}\cap U_{j})\rightarrow\cdots.$
We let $\mathfrak{U}_{i}$ be the formal schemes obtained by taking Witt vectors of this affine covering, with adic generic fibers $\mathfrak{U}_{i,C}$ over $C$. These form an affinoid perfectoid covering of the adic space $\mathfrak{Ig}^{b}_{C}$, and, by the acyclicity of affinoid perfectoids [Schperf, Theorem 1.8 (iv)], we have that
(8) ${}0\rightarrow\mathcal{O}(\mathfrak{Ig}_{C}^{b})\rightarrow\prod_{i\in I}\mathcal{O}(\mathfrak{U}_{i,C})\rightarrow\prod_{i,j\in I}\mathcal{O}(\mathfrak{U}_{i,C}\cap\mathfrak{U}_{j,C})\rightarrow\cdots.$
It follows that the Čech complex (8) is obtained from Čech complex (7) by taking Witt vectors followed by taking the completed tensor product with $C$. Using this, we deduce that the identification (6) follows from the identification (5), as desired.
Now $\mathcal{J}_{b}$ acts on the LHS of (6), as discussed above. Using that $\mathfrak{Ig}^{b,*}$ is affinoid, this will give the desired action on $\mathfrak{Ig}^{b,*}$, and one can see that it preserves the boundary by using the moduli interpretation of the toroidal compactifications, and the description of the $J_{b}(\mathbb{Q}_{p})$ action on cusp labels, as discussed in §2.2.2. ∎
@@
Lastly, we will consider the map $\overline{\pi}^{b}_{\mathfrak{Ig}}:[\overline{\mathfrak{Ig}}^{b}_{C}/\mathcal{J}_{b}]\rightarrow[\ast/\mathcal{J}_{b}]$, given by taking the canonical compactification of $\pi^{b}_{\mathfrak{Ig}}$, where we note, by [CS1, Proposition 4.2.22], the $v$-stack $[\mathop{\rm Spd}(C)/\mathcal{J}_{b}]$ is partially proper over $\mathop{\rm Spd}(C)$. We now have the following Proposition.
@@
(9) ${\left[(\mathcal{S}^{b,\circ,\mathrm{rd}}_{K^{p},C}\times_{\mathcal{F}\ell_{G,\mu^{-1}}}\mathcal{F}\ell_{G,\mu^{-1}}^{b})/\underline{G(\mathbb{Q}_{p})}\right]}$${\left[\mathcal{F}\ell^{b}_{G,\mu^{-1}}/\underline{G(\mathbb{Q}_{p})}\right]}$${\left[\mathfrak{Ig}_{C}^{b}/\mathcal{J}_{b}\right]}$${\left[\mathop{\rm Spd}(C)/\mathcal{J}_{b}\right]}$$\scriptstyle{\pi_{\mathrm{HT}}^{b,\circ}}$$\scriptstyle{\tilde{h}_{b}^{\leftarrow}}$$\scriptstyle{h^{\leftarrow}_{b}}$$\scriptstyle{\pi_{\mathfrak{Ig}}^{b}}$
@@
(10) ${\left[\overline{\mathcal{S}}^{b,\circ}_{K^{p},C}/\underline{G(\mathbb{Q}_{p})}\right]}$${\left[\mathcal{F}\ell^{b}_{G,\mu^{-1}}/\underline{G(\mathbb{Q}_{p})}\right]}$${\left[\overline{\mathfrak{Ig}}_{C}^{b}/\mathcal{J}_{b}\right]}$${\left[\mathop{\rm Spd}(C)/\mathcal{J}_{b}\right].}$$\scriptstyle{\overline{\pi}_{\mathrm{HT}}^{b,\circ}}$$\scriptstyle{\tilde{h}_{b}^{\leftarrow}}$$\scriptstyle{h^{\leftarrow}_{b}}$$\scriptstyle{\overline{\pi}_{\mathfrak{Ig}}^{b}}$
@@
Consider the moduli space of local shtukas $\mathrm{Sht}(G,b,\mu)_{\infty,C}$, as defined in [SW, §23]. This represents the functor sending $S\in\mathrm{Perf}_{C}$ to the set of all pairs $(S^{\#},\alpha)$ where $S^{\#}$ is the untilt of $S$ coming from the map $S\rightarrow\mathop{\rm Spd}(C)$, and $\alpha$ is a modification from $\mathcal{E}_{b}\dashrightarrow\mathcal{E}_{0}$ with meromorphy along $S^{\#}$ and bounded by $\mu$. We have a local Hodge-Tate period morphism
$\mathrm{Sht}(G,b,\mu)_{\infty,C}\rightarrow\mathcal{F}\ell_{G,\mu^{-1}}^{b},$
@@
(11) ${\mathrm{Sht}(G,b,\mu)_{\infty,C}}$${\mathop{\rm Spd}(C)}$${\mathcal{F}\ell_{G,\mu^{-1}}^{b}}$${\left[\mathop{\rm Spd}(C)/\mathcal{J}_{b}\right]}$
@@
$\mathcal{F}\ell^{b}_{G,\mu^{-1}}\times_{[\mathop{\rm Spd}(C)/\mathcal{J}_{b}]}[\mathfrak{Ig}_{C}^{b}/\mathcal{J}_{b}]\simeq\mathrm{Sht}(G,b,\mu)_{\infty,C}\times^{\mathcal{J}_{b}}\mathfrak{Ig}_{C}^{b}.$
Moreover, we see, by [CS1, Corollary 4.3.19,Lemma 4.3.20], that we have an isomorphism
(12) $\mathrm{Sht}(G,b,\mu)_{\infty,C}\times_{C}\mathfrak{Ig}_{C}^{b}\simeq\mathcal{S}^{b,\circ,\mathrm{rd}}_{K^{p}}\times_{\mathcal{F}\ell_{G,\mu^{-1}}}\mathrm{Sht}(G,b,\mu)_{\infty,C},$
and again applying (11) implies that $\mathcal{S}^{b,\circ,\mathrm{rd}}_{K^{p}}\times_{\mathcal{F}\ell_{G,\mu^{-1}}}\mathcal{F}\ell_{G,\mu^{-1}}^{b}$ is isomorphic to the quotient of $\mathcal{S}^{b,\circ,\mathrm{rd}}_{K^{p}}\times_{\mathcal{F}\ell_{G,\mu^{-1}}}\mathrm{Sht}(G,b,\mu)_{\infty,C}$ by the action of $\mathcal{J}_{b}$ (here $\mathcal{J}_{b}$ acts via the action on the second factor). Thus, we have an isomorphism:
$\mathrm{Sht}(G,b,\mu)_{\infty,C}\times^{\mathcal{J}_{b}}\mathfrak{Ig}_{C}^{b}\simeq\mathcal{S}^{b,\circ,\mathrm{rd}}_{K^{p}}\times_{\mathcal{F}\ell_{G,\mu^{-1}}}\mathcal{F}\ell_{G,\mu^{-1}}^{b}.$

### v5

@@ 3.2. Mantovan’s Product Formula
The diamond attached to the perfectoid Igusa variety $\mathfrak{Ig}^{b}_{C}$ comes equipped with an action of $\mathcal{J}_{b}$. In particular, we consider the formal group scheme $\mathrm{Aut}_{G}(\tilde{\mathbb{X}}_{b})$ over $\mathrm{Spf}(W(\overline{\mathbb{F}}_{p}))$ described in [7, Definition 4.2.9] parametrizing automorphisms of the universal cover of $\mathbb{X}_{b}$. We write $\mathrm{Aut}_{G}(\tilde{\mathbb{X}}_{b})_{\overline{\mathbb{F}}_{p}}$ for the special fiber of this formal group scheme, which appeared in 2.2.1. This defines a formal group over $\overline{\mathbb{F}}_{p}$, which we regard as an fpqc sheaf on schemes over $\overline{\mathbb{F}}_{p}$. The formal group scheme $\mathrm{Aut}_{G}(\tilde{\mathbb{X}}_{b})$ over $\mathrm{Spf}(W(\overline{\mathbb{F}}_{p}))$ is then given by taking Witt vectors of this formal group.
As is done in [7, Corollary 4.3.5], one can, using the moduli description of $\mathrm{Ig}^{b}$ provided in (3), show that the fpqc sheaf $\mathrm{Aut}_{G}(\tilde{\mathbb{X}}_{b})_{\overline{\mathbb{F}}_{p}}$ acts on $\mathrm{Ig}^{b}$, by noting that quasi-isogenies of $\mathbb{X}_{b}$ are the same as automorphisms of the universal cover.
Functorality of Witt vectors then gives rise to an action of $\mathrm{Aut}_{G}(\tilde{\mathbb{X}}_{b})$ on the Witt vector lift of $\mathrm{Ig}^{b}$, and passing to rigid generic fibers over $C$ gives an action of the rigid generic fiber of $\mathrm{Aut}_{G}(\tilde{\mathbb{X}}_{b})$ on $\mathfrak{Ig}^{b}_{C}$, which we denote by $\mathrm{Aut}_{G}(\tilde{\mathbb{X}}_{b})_{C}$. By [7, Lemma 4.2.10], $\mathrm{Aut}_{G}(\tilde{\mathbb{X}}_{b})_{C}$ is a perfectoid space which we identify with its associated $v$-sheaf on $\mathrm{Perf}_{C}$ using the Yoneda embedding. The following seems to be well known, but we sketch a proof as we could not find it explicitly written down in the literature.
Proposition 3.8.
There is an identification
(9) $\mathrm{Aut}_{G}(\tilde{\mathbb{X}}_{b})_{C}\simeq\mathcal{J}_{b}$
of group-valued functors on $\mathrm{Perf}_{C}$, where $\mathcal{J}_{b}$ is the group diamond parametrizing automorphisms of the $G$-bundle $\mathcal{E}_{b}$ on the Fargues-Fontaine curve attached to $b$.
Proof.
(Sketch) We fix $(\mathbb{X}_{b})_{O_{C}}$ to be a lift of $\mathbb{X}_{b}$ up to quasi-isogeny (with its $G$-structure). For an object $S=\mathop{\rm Spa}(R,R^{+})\in\mathrm{Perf}_{C}$, the $S$-points of LHS identifies with the set of formal quasi-isogenies
$(\mathbb{X}_{b})_{O_{C}}\times_{\mathrm{Spf}(O_{C})}\mathrm{Spf}(R^{+})\dashrightarrow(\mathbb{X}_{b})_{O_{C}}\times_{\mathrm{Spf}(O_{C})}\mathrm{Spf}(R^{+})$
in the sense of [10, Section 2.5], respecting the $G$-structure. In other words, this is a quasi-isogeny
$(\mathbb{X}_{b})_{O_{C}}\times_{O_{C}}R^{+\flat}/\varpi\dashrightarrow(\mathbb{X}_{b})_{O_{C}}\times_{O_{C}}R^{+\flat}/\varpi$
respecting the additional structures, where $\varpi\in R^{+\flat}$ is a suitable choice of pseudo-uniformizer in the tilt $R^{+\flat}$ of $R^{+}$, and the quotient map is given by the map $R^{+}\rightarrow R^{+\flat}/\varpi$ induced by Fontaine’s map $\theta_{(R,R^{+})}:W(R^{+\flat})\rightarrow R^{+}$, and the quasi-isogeny doesn’t depend on the choice of pseudo-uniformizer by Serre-Tate theory.
As described in [10, Section 2.5.3], to a $p$-divisible $G_{0}$ group over $R^{+}/\varpi$, one can associated a vector bundle $\mathcal{E}(G_{0})$ on $X_{S}$ using crystalline Dieudonné theory. As explained in [10, Section 2.5.6], a quasi-isogeny $G_{0}\dashrightarrow G_{0}$ induces an isomorphism of vector bundles
$\mathcal{E}(G_{0})\xrightarrow{\simeq}\mathcal{E}(G_{0})$
on $X_{S}$. By applying the Tannakian-formalism, it follows that for a $p$-divisible group $G_{0}$ with $G$-structure and a quasi-isogeny $G_{0}\dashrightarrow G_{0}$ respecting the additional structures, that one obtains a $G$-bundle $\mathcal{E}_{G}(G_{0})$ and an isomorphism of $G$-bundles $\mathcal{E}_{G}(G_{0})\xrightarrow{\simeq}\mathcal{E}_{G}(G_{0})$. Applying this to $G_{0}=(\mathbb{X}_{b})_{O_{C}}\times_{O_{C}}R^{+\flat}/\varpi$, one can check by our assumption that $(\mathbb{X}_{b})_{O_{C}}$ was a lift of $\mathbb{X}_{b}$ and [16, Theorem II.0.4], that one has pro-étale locally an identification $\mathcal{E}_{G}(G_{0})=\mathcal{E}_{b,S}$, where $\mathcal{E}_{b,S}$ is the pullback of $\mathcal{E}_{b}$ the vector bundle attached to the isocrystal $b\in B(G)$ on $X_{C}$ along the natural map $X_{S}\rightarrow X_{C}$. This defines the desired natural transformation from the functor of points of the LHS to the functor of points of the RHS of (9). In order to check it is an equivalence, we use that the LHS and RHS of (9) are partially proper over $\mathop{\rm Spa}(C,O_{C})$ to reduce to checking it on $(R,R^{\circ})$-points. For the RHS, this follows from the definition, as the category of vector bundles on $X_{(R,R^{+})}$ and $X_{(R,R^{\circ})}$ are equivalent, and for the LHS it follows from [7, Proposition 4.2.22]. The desired equivalence is then reduced to the fully faithfullness statement in [50, Proposition 2.2.7], via [10, Lemma 2.5.5].
Using the previous Proposition, one sees that the action of $\mathrm{Aut}_{G}(\tilde{\mathbb{X}}_{b})_{C}$ on $\mathfrak{Ig}^{b}_{C}$ described above gives rise to an action of the group diamond $\mathcal{J}_{b}$ parametrizing automorphism of $\mathcal{E}_{b}$.
This allows us to form the $v$-stack quotient $\mathfrak{Ig}_{C}^{b}/\mathcal{J}_{b}]$, which is equipped with a natural map:
@@
We would like to say $\pi_{\mathfrak{Ig}}^{b}$ pulls back to the map $\pi^{b}_{\mathrm{HT}}$ along the map (8). However, as seen in Corollary 3.5, we need to account for the additional points in the fiber of the Hodge-Tate period morphism that are not seen by the perfectoid Igusa varieties $\mathfrak{Ig}^{b}_{C}$. To capture this, we need to show that $\pi_{\mathfrak{Ig}}^{b}$ also extends to the partial minimal compactification in a suitable sense. We recall that we have a decomposition
(10) $\mathcal{J}_{b}\simeq\underline{J_{b}(\mathbb{Q}_{p})}\ltimes\mathcal{J}_{b}^{>0},$
where $\underline{J_{b}(\mathbb{Q}_{p})}$ is the constant group scheme coming from the automorphisms of $\mathcal{E}_{b}$ which preserve the graded pieces of the Harder-Narasimhan filtration and $\mathcal{J}_{b}^{>0}$ is an $\ell$-adically contractible unipotent group diamond of dimension $\langle 2\rho_{G},\nu_{b}\rangle$. Analogously, we have a decomposition of fpqc sheaves
(11) $\mathrm{Aut}_{G}(\tilde{\mathbb{X}}_{b})_{\overline{\mathbb{F}}_{p}}=\underline{J_{b}(\mathbb{Q}_{p})}\ltimes\widetilde{\mathcal{U}}_{b},$
where $J_{b}(\mathbb{Q}_{p})$ can be identified with the group of self-quasi-isogenies of $\mathbb{X}_{b}$ over $\overline{\mathbb{F}}_{p}$ that respect the $G$-structure, $\underline{J_{b}(\mathbb{Q}_{p})}$ is the associated constant group scheme, and $\tilde{\mathcal{U}_{b}}$ is a unipotent formal group scheme, see [7, Proposition 4.2.11]. We now have the following.
@@
Assuming 1.13, the action of $\mathcal{J}_{b}$ on the perfectoid Igusa variety $\mathfrak{Ig}^{b}_{C}$ extends uniquely to an action on $\mathfrak{Ig}^{b,*}_{C}$. In particular, by functorality of the formation of the canonical compactification ([56, Proposition 18.6]), we have a map
@@
extending $\pi_{\mathfrak{Ig}}^{b}$. This restriction of the action to the subgroup $\underline{J_{b}(\mathbb{Q}_{p})}\subset\mathcal{J}_{b}$ also preserves the boundary $\partial\mathfrak{Ig}^{b,*}$, so, in particular, we also get a map
$\tilde{\pi}_{\mathfrak{Ig}}^{b,\partial}:[(\overline{\mathfrak{Ig}}^{b,*}_{C}\setminus\overline{\partial\mathfrak{Ig}}^{b,*}_{C})/\underline{J_{b}(\mathbb{Q}_{p})}]\rightarrow[\mathop{\rm Spd}(C)/\underline{J_{b}(\mathbb{Q}_{p})}]$
by pulling back $\pi_{\mathfrak{Ig}}^{b,*}$ along the map $\mathop{\rm Spd}(C)/\underline{J_{b}(\mathbb{Q}_{p})}]\rightarrow[\mathop{\rm Spd}(C)/\mathcal{J}_{b}]$ induced by the decomposition (10) and restricting.
@@
as explained in the proof of Proposition 3.6. Under Assumption 1.13, we can apply the algebraic form of Hartogs’ principle (See for example [57, Proposition III.2.9]) to the open inclusion $g_{b,m}$ to conclude an isomorphism of global sections via restriction. This gives the corresponding claim for the map $g_{b}$ of perfect schemes, by taking inverse limits and perfectifying. In particular, we have an isomorphism
@@
on global sections. Since the perfect scheme $\mathrm{Ig}^{b,*}$ is affine, the isomorphism of global sections implies that the action of the formal group scheme $\mathrm{Aut}_{G}(\tilde{\mathbb{X}}_{b})$ regarded as a fpqc sheaf on the perfect scheme $\mathrm{Ig}^{b}$ extends to an action on the perfect affine scheme $\mathrm{Ig}^{b,*}$. As before, by taking Witt vectors and passing to the rigid generic fiber over $C$, this induces an action of $\mathrm{Aut}_{G}(\tilde{\mathbb{X}}_{b})_{C}$ on the perfectoid space $\mathfrak{Ig}^{b,*}_{C},$ extending the action on $\mathfrak{Ig}^{b}_{C}$. Passing through the identification $\mathrm{Aut}_{G}(\tilde{\mathbb{X}}_{b})_{C}\simeq\mathcal{J}_{b}$ of Proposition 3.8, this gives us the desired action of $\mathcal{J}_{b}$ on $\mathfrak{Ig}^{b,*}_{C}$.
Now, to understand the preservation of the boundary, we understand the construction of the action of $\mathcal{J}_{b}$ in a more abstract way. In particular, we consider the small diamond functor $(-)^{\diamond}$ of [22, Section 3.1]. Recall $\mathrm{Perf}$ is the category of perfectoid spaces in characteristic $p$ over $\overline{\mathbb{F}}_{p}$, the $v$-sheaf $\mathrm{Aut}_{G}(\tilde{\mathbb{X}}_{b})^{\diamond}$ on $\mathrm{Perf}$ attached to the formal group $\mathrm{Aut}_{G}(\tilde{\mathbb{X}}_{b})^{\diamond}$ is defined by the functor sending $(R,R^{+})\in\mathrm{Perf}$ to the group
[…]
for some pseudo-uniformizer $\varpi$, where the last equivalence follows from Serre-Tate lifting of $p$-divisible groups along $p$-nilpotent rings, as in [8, Theorem 2.4.1]. The $v$-sheaves $\mathrm{Ig}^{b,\diamond}$ (resp. $\mathrm{Ig}^{b,*,\diamond}$) are defined by sending $(R,R^{+})$ to $\mathrm{Ig}^{b}(\mathop{\rm Spec}R^{+})$ and $\mathrm{Ig}^{b,*}(\mathop{\rm Spec}R^{+})$. We note that, since $\mathrm{Ig}^{b}$ and $\mathrm{Ig}^{b,*}$ are perfect and $R^{+}$ is perfectoid, we have identifications $\mathrm{Ig}^{b}(\mathop{\rm Spec}R^{+})\simeq\mathrm{Ig}^{b}(\mathop{\rm Spec}R^{+}/\varpi)$ and $\mathrm{Ig}^{b,*}(\mathop{\rm Spec}R^{+})\simeq\mathrm{Ig}^{b,*}(\mathop{\rm Spec}R^{+}/\varpi)$. In particular, we may understand the action of $\mathrm{Aut}_{G}(\tilde{\mathbb{X}}_{b})^{\diamond}$ on the $v$-sheaves $\mathrm{Ig}^{b,\diamond}$ and $\mathrm{Ig}^{b,*,\diamond}$ in terms of the action of $\mathrm{Aut}_{G}(\tilde{\mathbb{X}}_{b})(\mathop{\rm Spec}R^{+}/\varpi)$ on $\mathrm{Ig}^{b}(\mathop{\rm Spec}R^{+}/\varpi)$ and $\mathrm{Ig}^{b,*}(\mathop{\rm Spec}R^{+}/\varpi)$.
The action of $\mathcal{J}_{b}$ on $\mathfrak{Ig}^{b}_{C}$ and $\mathfrak{Ig}^{b,*}_{C}$ described above may now be given by base-changing $\mathrm{Aut}_{G}(\tilde{\mathbb{X}}_{b})^{\diamond}$, $\mathrm{Ig}^{b}$, $\mathrm{Ig}^{b,*}_{C}$ and base-changing along $\mathop{\rm Spd}(C)\rightarrow\mathop{\rm Spd}(\overline{\mathbb{F}}_{p})$ in the category of $v$-sheaves and identifying the resulting functors on $\mathrm{Perf}_{C}$ with $\mathcal{J}_{b}\simeq\mathrm{Aut}_{G}(\tilde{\mathbb{X}}_{b})_{C}$, $\mathfrak{Ig}^{b}_{C}$, and $\mathfrak{Ig}^{b,*}_{C}$. In particular, the relationship between the construction described above in terms of taking Witt vectors and the passing to the rigid generic fiber over $C$ follows from from the relationship between untilts of a characteristic $p$ perfectoid $(R,R^{+})$ and Cartier divisors in the Witt ring $W(R^{+})$, as described in [55, Lemma 6.28].
We now use this perspective to show that the action of sub group-sheaf $\underline{J_{b}(\mathbb{Q}_{p})}\subset\mathrm{Aut}_{G}(\tilde{\mathbb{X}}_{b})^{\diamond}$ given by the decomposition (11) preserves the closed (by the valuative criterion of properness) subsheaf $\partial\mathrm{Ig}^{b,*,\diamond}\subset\mathrm{Ig}^{b,*,\diamond}$. By base-changing along $\mathop{\rm Spd}(C)\rightarrow\mathop{\rm Spd}(\overline{\mathbb{F}}_{p})$, this will give the desired claim.
[…]
Lemma 3.10.
Suppose we have a group-valued fpqc sheaf $G/\overline{\mathbb{F}}_{p}$ acting on a scheme $X/\overline{\mathbb{F}}_{p}$ via $m:G\times X\rightarrow X$ and suppose that $Z\subset X$ is a closed subscheme which is reduced. We write $U:=X\setminus Z$ for the open complement and assume the action of $G$ preserves $U\subset X$. Then, for all $T/\overline{\mathbb{F}}_{p}$ reduced, the action of the $T$-points $G(T)$ on $X(T)$ preserves $Z(T)$.
Proof.
For $T$ as in the statement, consider a $T$-point $x:T\rightarrow Z\rightarrow X$ and an element $g\in G(T)$. We need to show that $m(g,x):T\rightarrow X$ factors through $Z\rightarrow X$. Since $T$ and $Z$ are reduced, it suffices by [61, Lemma 0356], to show that the induced map $|m(g,x)|:|T|\rightarrow|X|$ on the underlying topological space factors through $|Z|\rightarrow|X|$. Suppose for the sake of contradiction that it does not, then there exists a geometric point $\mathop{\rm Spec}(k)\rightarrow T$ factoring through $U\rightarrow X$. Now if we consider $m(g^{-1},m(g,x))=x:T\rightarrow X$, then we arrive at a contradiction since the precomposition of this map with $\mathop{\rm Spec}(k)\rightarrow T$ must factor through $U\subset X$, by the assumption that $G$ preserves $U$. This however contradicts the decomposition $|X|=|U|\sqcup|Z|$ on topological spaces. ∎
Remark 3.11.
It is true that the full action of $\mathcal{J}_{b}$ preserves the Zariski closed boundary $\partial\mathfrak{Ig}^{b,*}_{C}\subset\mathfrak{Ig}^{b,*}_{C}$. This is shown in [6, Section 4-5] (In fact, a stronger claim is shown that $\mathcal{J}_{b}$ preserves all the closed boundary strata of $\partial\mathfrak{Ig}^{b,*}_{C}$). However, this requires a much more detailed argument involving lifting to the toroidal compactification. As a result, we will not address this here.
@@
Lastly, we will consider the map $\overline{\pi}^{b}_{\mathfrak{Ig}}:[\overline{\mathfrak{Ig}}^{b}_{C}/\mathcal{J}_{b}]\rightarrow[\mathop{\rm Spd}(C)/\mathcal{J}_{b}]$, given by taking the canonical compactification of $\pi^{b}_{\mathfrak{Ig}}$, where we note, by [7, Proposition 4.2.22], the $v$-stack $\mathop{\rm Spd}(C)/\mathcal{J}_{b}]$ is partially proper over $\mathop{\rm Spd}(C)$. We now have the following Proposition.
@@
(13) $\left[(\mathcal{S}^{b,\circ,\mathrm{rd}}_{K^{p},C}\times_{\mathcal{F}\ell^{\Diamond}_{G,\mu^{-1}}}\mathcal{F}\ell_{G,\mu^{-1}}^{b,\Diamond})/\underline{G(\mathbb{Q}_{p})}\right]$$\left[\mathcal{F}\ell^{b,\Diamond}_{G,\mu^{-1}}/\underline{G(\mathbb{Q}_{p})}\right]$$\left[\mathfrak{Ig}_{C}^{b}/\mathcal{J}_{b}\right]$$\left[\mathop{\rm Spd}(C)/\mathcal{J}_{b}\right]$$\scriptstyle{\pi_{\mathrm{HT}}^{b,\circ}}$$\scriptstyle{\tilde{h}_{b}^{\leftarrow}}$$\scriptstyle{h^{\leftarrow}_{b}}$$\scriptstyle{\pi_{\mathfrak{Ig}}^{b}}$
@@
(14) $\left[\overline{\mathcal{S}}^{b,\circ}_{K^{p},C}/\underline{G(\mathbb{Q}_{p})}\right]$$\left[\mathcal{F}\ell^{b,\Diamond}_{G,\mu^{-1}}/\underline{G(\mathbb{Q}_{p})}\right]$$\left[\overline{\mathfrak{Ig}}_{C}^{b}/\mathcal{J}_{b}\right]$$\left[\mathop{\rm Spd}(C)/\mathcal{J}_{b}\right].$$\scriptstyle{\overline{\pi}_{\mathrm{HT}}^{b,\circ}}$$\scriptstyle{\tilde{h}_{b}^{\leftarrow}}$$\scriptstyle{h^{\leftarrow}_{b}}$$\scriptstyle{\overline{\pi}_{\mathfrak{Ig}}^{b}}$
@@
Consider the moduli space of local shtukas $\mathrm{Sht}(G,b,\mu)_{\infty,C}$, as defined in [55, Definition 23.1.1]. This represents the functor sending $S\in\mathrm{Perf}_{C}$ to the set of all pairs $(S^{\#},\alpha)$ where $S^{\#}$ is the untilt of $S$ coming from the map $S\rightarrow\mathop{\rm Spd}(C)$, and $\alpha$ is a modification from $\mathcal{E}_{b}\dashrightarrow\mathcal{E}_{0}$ with meromorphy along $S^{\#}$ and bounded by $\mu$. We have a local Hodge-Tate period morphism
$\mathrm{Sht}(G,b,\mu)_{\infty,C}\rightarrow\mathcal{F}\ell_{G,\mu^{-1}}^{b,\Diamond},$
@@
(15) $\mathrm{Sht}(G,b,\mu)_{\infty,C}$$\mathop{\rm Spd}(C)$$\mathcal{F}\ell_{G,\mu^{-1}}^{b,\Diamond}$$\left[\mathop{\rm Spd}(C)/\mathcal{J}_{b}\right]$}
@@
$\mathcal{F}\ell^{b,\Diamond}_{G,\mu^{-1}}\times_{[\mathop{\rm Spd}(C)/\mathcal{J}_{b}]}[\mathfrak{Ig}_{C}^{b}/\mathcal{J}_{b}]\simeq\mathrm{Sht}(G,b,\mu)_{\infty,C}\times^{\mathcal{J}_{b}}\mathfrak{Ig}_{C}^{b}.$
Moreover, we see, by [7, Corollary 4.3.19,Lemma 4.3.20], that we have an isomorphism
(16) $\mathrm{Sht}(G,b,\mu)_{\infty,C}\times_{C}\mathfrak{Ig}_{C}^{b}\simeq\mathcal{S}^{b,\circ,\mathrm{rd}}_{K^{p}}\times_{\mathcal{F}\ell^{\Diamond}_{G,\mu^{-1}}}\mathrm{Sht}(G,b,\mu)_{\infty,C}.$
Moreover, this map is $\mathcal{J}_{b}$-equivariant. Here $\mathcal{J}_{b}$ acts on the LHS via $(x,y)\mapsto(jx,j^{-1}y)$, and $\mathcal{J}_{b}$ acts on the RHS by the trivial action on $\mathcal{S}^{b,\circ,\mathrm{rd}}_{K^{p}}$ and the natural action on $\mathrm{Sht}(G,b,\mu)_{\infty,C}$ by automorphisms of bundles.
Now applying (15) implies that $\mathcal{S}^{b,\circ,\mathrm{rd}}_{K^{p}}\times_{\mathcal{F}\ell^{\Diamond}_{G,\mu^{-1}}}\mathcal{F}\ell_{G,\mu^{-1}}^{b,\Diamond}$ is isomorphic to the quotient of $\mathcal{S}^{b,\circ,\mathrm{rd}}_{K^{p}}\times_{\mathcal{F}\ell^{\Diamond}_{G,\mu^{-1}}}\mathrm{Sht}(G,b,\mu)_{\infty,C}$ by the action of $\mathcal{J}_{b}$ (here $\mathcal{J}_{b}$ acts via the action on the second factor). Thus, we have an isomorphism:
$\mathrm{Sht}(G,b,\mu)_{\infty,C}\times^{\mathcal{J}_{b}}\mathfrak{Ig}_{C}^{b}\simeq\mathcal{S}^{b,\circ,\mathrm{rd}}_{K^{p}}\times_{\mathcal{F}\ell^{\Diamond}_{G,\mu^{-1}}}\mathcal{F}\ell_{G,\mu^{-1}}^{b,\Diamond}.$